# Locating and Editing Factual Associations in Mamba

### v1

@@ Abstract
We investigate the mechanisms of factual recall in the Mamba state space model. Our work is inspired by previous findings in autoregressive transformer language models suggesting that their knowledge recall is localized to particular modules at specific token locations; we therefore ask whether factual recall in Mamba can be similarly localized. To investigate this, we conduct four lines of experiments on Mamba. First, we apply causal tracing or interchange interventions to localize key components inside Mamba that are responsible for recalling facts, revealing that specific components within middle layers show strong causal effects at the last token of the subject, while the causal effect of intervening on later layers is most pronounced at the last token of the prompt, matching previous findings on autoregressive transformers. Second, we show that rank-one model editing methods can successfully insert facts at specific locations, again resembling findings on transformer models. Third, we examine the linearity of Mamba’s representations of factual relations. Finally we adapt attention-knockout techniques to Mamba to dissect information flow during factual recall. We compare Mamba directly to a similar-sized transformer and conclude that despite significant differences in architectural approach, when it comes to factual recall, the two architectures share many similarities.
@@ 1 Introduction
Studies of autoregressive transformer language models’ (LMs) processing of factual statements such as The Eiffel Tower is located in Paris, have identified a localized pattern of internal computations when recalling facts (Meng et al., 2022a; b; Geva et al., 2023; Hernandez et al., 2023; Nanda et al., 2023), and have further found that those LMs can be edited by making single-layer rank-one changes in model parameters to alter a specific fact. Although these localized phenomena appear to generalize across transformer LMs, the extent to which similar locality might appear in very different architectures—such as recurrent networks (RNNs)—has not yet been investigated.
@@
Our paper is a case study confronting a key methodological challenge that broadly faces interpretability researchers: as state-of-the-art neural network architectures evolve, we must ask, can the detailed analytical methods and tools developed for one neural architecture, such as transformer LMs, be generalized and applied to a different neural architecture such as Mamba? In this we are able to answer the question with a qualified “yes”: we find that many of the methods used to analyze transformers can also provide insights on Mamba. We also discuss mismatches—that is, interpretation methods (such as path-dependent attention patching) that do not transfer to Mamba as easily due to architectural differences.
@@
Finally, we apply methods for understanding the overall information flows in Mamba. Inspired by the linearity findings of Hernandez et al. (2023), we measure the linearity of the relations between subject and object embeddings. And inspired by Geva et al. (2023), we examine information flow by adapting attention-blocking methods to the attention-free Mamba architecture.
@@ 2 Background on Mamba
Mamba, introduced in Gu & Dao (2023), is a recent family of language models based on state space models (SSMs). SSMs are designed to model the evolution of hidden state across time with a first-order differential equation (Koopman et al., 1999; Durbin & Koopman, 2012), and when they are used as the recurrent state of an RNN, they can enable highly efficient parallelized training (Gu et al., 2021). To achieve good performance in language modeling, the Mamba SSM introduces input-dependent parameterization or selective-SSM instead of the traditional time-invariant SSMs. Mamba uses a special architecture called MambaBlock [^1] In their paper, Gu & Dao (2023) call this component Mamba—the same name as the LM family., which is stacked homogeneously, replacing both attention and MLP blocks used in transformer layers. Here, we focus on the different operations performed inside a MambaBlock.
@@ 3 Locating Key States for Factual Recall
Figure 2: (a) Activation patching. A state from the clean run $G$ is patched into its corresponding position in $G^{*}$. This has a downstream effect of potentially changing all the states that depend on the patched state in $G^{*}[\leftarrow h_{i}^{(\ell)}]$. (b) Average indirect effect of applying causal tracing on residual stream states ($h_{i}^{(\ell)}$ in Figure 1) across 400 different facts.
@@
This subject-swapping approach follows the recommendation of Zhang & Nanda (2023) and has the advantage of using a natural text perturbations to avoid introducing out-of-domain states to the model’s computation, as may happen with corrupting $s$ embeddings with Gaussian noise (used in Meng et al. (2022a)).
@@
Let $p(o)$, $p^{*}(o)$, and $p^{*}[{\leftarrow h_{i}^{(\ell)}}](o)$ denote the probability assigned to the correct answer $o$ in $G$, $G^{*}$, and $G^{*}[{\leftarrow h_{i}^{(\ell)}}]$ respectively. To measure the contribution of $h_{i}^{(\ell)}$ in recalling the fact $(s,r,o)$, we define the indirect effect (IE) as:
@@
In Figure 2(b) we plot the average indirect effect of restoring the residual states $h_{i}^{(\ell)}$ across different layer-token positions over 400 facts. The high IE in later layers at the last token position is natural, as restoring a clean $h_{i}^{(\ell)}$ there means restoring most of the model computation from $G$. However, Mamba also shows high causality in the middle layers at the last subject token position. This is consistent with what Meng et al. (2022a) observed in GPT family of language models.
@@
Figure 3: Average indirect effect of different states $o_{i}^{(\ell)}$, $g_{i}^{(\ell)}$, and $s_{i}^{(\ell)}$ over 400 facts. For each layer $\ell$, states for a window of 10 layers around $\ell$ are restored from the clean run $G$.
@@
Figure 4: A probe for path-specific effects. $h_{i}^{(\ell)}$ is restored from the clean run $G$ as in Figure 2(a). Then, to reveal the role of the SSM component, $s_{i}$ states from the corrupted run $G^{*}$ are also patched to block the prevent contributions on the $s_{i}$ paths.
@@
In Figure 3 we plot the average IE for $o_{i}^{(\ell)}$, $g_{i}^{(\ell)}$, and $s_{i}^{(\ell)}$. The plot for $o_{i}^{(\ell)}$ (Figure 3a) looks very similar to Figure 2(b), confirming that the output from MambaBlock has strong causal effects at both early and late sites. Interestingly, Figure 3c shows that the selective-SSM output $s_{i}^{(\ell)}$ has high IE in later layers at the last token of the prompt, resembling the behavior of attention modules in GPT models (Meng et al., 2022a). However, there is no state that appears to do the opposite; in other words, there is no state with strong effects at the early site and not at the late site. (The gate output $g_{i}^{(\ell)}$ does have stronger IE at the early site than the late site, but the effects are very weak.) A comparison with the comparable Pythia-2.8b transformer LMs is shown in Figure 9 in Appendix C. This comparison reveals a key way that Mamba differs from transformers: while transformer MLP outputs have effects in the early site and not the late site, in Mamba there is no similar state that specializes only at the early site at which factual recall would be expected to occur. This presents the question: which parameters in Mamba mediate factual recall?
@@
To investigate this question, we replicate an experiment from Meng et al. (2022a) to probe path-specific effects (Pearl, 2022) for paths that avoid either $g_{i}$, $s_{i}$, or $o_{i}$ contributions. First, in the corrupted run $G^{*}$, at token position $i$, we cache all the contributions from the $s_{i}$ paths as $s_{i}^{*}=\{s_{i}^{*(\ell)}|\;\ell\in[1,L]\}$. Then in the patched run $G^{*}[{\leftarrow h_{i}^{(\ell)}}]$ we patch $h_{i}^{(\ell)}$ cached from the clean run $G$ into its corresponding state (as in Figure 2(a)), but with an additional modification: to understand the contribution from the $s_{i}$ paths, we sever those paths by patching $s_{i}^{*}$ to their corresponding locations (see Figure 4). The same experiment can be conducted to understand the contributions of $g_{i}$ and $o_{i}$. We note that blocking the $o_{i}$ paths will block the $s_{i}$ and $g_{i}$ paths as well (see Figure 1).
@@
In Figure 5 we plot the average results of this experiment done at token positions (a) $i=\textit{subject last}$ and (b) $i=\textit{prompt last}$ over 400 examples. The key findings can be understood by examining the gap between the purple bars and the green, red, and blue bars: a large gap indicates a strong mediating role for SSM, $\text{W}_{g}$, or $\text{W}_{o}$ parameters, respectively. At the early site at the subject last token, both SSM and $\text{W}_{g}$ have a strong role, but $\text{W}_{o}$ plays an even larger role. Yet the strongest mediator at the late site is also $\text{W}_{o}$. This experiment highlights the importance of $\text{W}_{o}$ in both stages of predicting a fact, but it also suggests that Mamba does not separate early-site factual recall between these groups of parameters as cleanly as transformers. However, Figure 5 reveals a clean separation of roles between early to mid and later layers, similar to what Hernandez et al. (2023) observed in transformer LMs. However, we note that this division of responsibilities between layers can be more sharply noticed in Mamba when compared to transformers LMs (compare with Figure 10).
@@ 4 Editing Facts With ROME
Having begun to characterize the locations of important states for factual recall, we now investigate whether factual recall behavior can be edited. In particular, we apply the ROME (Rank One Model Editing, Meng et al., 2022a) technique to Mamba. ROME begins with the observation that any linear transformations can be considered as an associative memory (Anderson, 1972; Kohonen, 1972), mapping a set of keys $K=[k_{1}|k_{2}|\dots]$ to their corresponding values $V=[v_{1}|v_{2}|\dots]$, and uses this to edit factual associations in transformer LMs. Here, we apply the technique to a particular set of linear transformations within Mamba, and report our editing success on each. [^2] Further motivating these experiments, previous work has shown that the locations identified by activation patching techniques are often not those which have the strongest edit performance  (Hase et al., 2024).
The input to ROME is a prompt $x=(s,r)$, where $s$ (Emmanuel Macron) is a subject entity and $r$ (is the President of) is a relation. ROME also takes a counterfactual object $o^{*}$ (England), meant to replace the correct object $o$ (France) in the model’s output. To effect that change, ROME generates a rank-one update to $\text{W}_{down}^{(\ell)}$, the down-projection matrix of the MLP module for the last token of the subject at layer $\ell$—which plays the role of the associative memory. In generating the rank-one update, ROME considers the input to $\text{W}_{down}^{(\ell)}$ as the key ($k_{*}$). Then, with gradient descent ROME calculates a value ($v_{*}$) such that, when $v_{*}$ is inserted as the output of $\text{W}_{down}^{(\ell)}$, the model will output $o^{*}$. Importantly, while optimizing $v_{*}$, ROME attempts to minimize unrelated changes in model outputs (Joe Biden, for example, should still be mapped to the United States post-edit). Finally, ROME adds a rank-1 matrix $\Delta$ to $\text{W}_{down}^{(\ell)}$ such that $\big{(}\text{W}_{down}^{(\ell)}+\Delta\big{)}k_{*}\approx v_{*}$. (For mathematical and implementation details see Meng et al. (2022a).)
@@ 4.1 Applying ROME in Mamba
We apply ROME to three different projection matrices: $\text{W}_{a}^{(\ell)}$ which affects only the selective-SSM path, $\text{W}_{g}^{(\ell)}$ which affects only the gating path, and $\text{W}_{o}^{(\ell)}$, the final output of the MambaBlock, which is added to the residual. We plot ROME performance across different layers and projection matrices ($\text{W}_{a}^{(\ell)}$, $\text{W}_{g}^{(\ell)}$, and $\text{W}_{o}^{(\ell)}$) in Figure 6a.
@@
To evaluate editing performance, we use the CounterFact dataset from Meng et al. (2022a). CounterFact contains 20K counterfactual examples in the form $(s,r,o\rightarrow o^{*})$, where $o$ is the correct answer to the prompt $x=(s,r)$, and $o^{*}$ is the object which is to be inserted as the new answer to the prompt. (See Section A.1 for details). We select the first 2000 examples from this dataset for our module-layer sweep. We use the original evaluation matrices in Meng et al. (2022a) to measure ROME edit performance.
@@
    Generalization (PS): A successful edit should be persistent across different paraphrases of $(s,r)$. For each of the request instances $(s,r,o\rightarrow o^{*})$, $p(o^{*})>p(o)$ is checked post-edit with a set of different rephrasings $x_{p}\sim\mathcal{P}(s,r)$ of the prompt $x=(s,r)$, where $\mathcal{P}$ denotes a set of paraphrased templates for the relation $r$.
@@
    Specificity (NS): Finally, the edit should be specific to $\mathcal{P}(s,r)$ and should not additionally change the mapping of some nearby subject $s_{n}$ to $o^{*}$. To evaluate the specificity of an edit we measure $p(o_{n})>p(o^{*})$ with $\mathcal{P}(s_{n},r)$ for a set of nearby factual associations $\{(s_{n},r,o_{n})\,|\,o_{n}\neq o^{*}\}$.
@@
Figure 6a shows that ROME can achieve high scores (S) for a range of early to middle layers by modifying any one of the projection matrices $\text{W}_{a}^{(\ell)}$, $\text{W}_{g}^{(\ell)}$, or $\text{W}_{o}^{(\ell)}$, matching observations made by Hase et al. (2024) regarding transformer LMs. However, performance does depend on the location of the edit. For example, in the case of $\text{W}_{g}^{(\ell)}$ and $\text{W}_{o}^{(\ell)}$, the score (S) and generalization (PS) drops after layer 43. This is consistent with our findings from the path-blocking experiment in Figure 5a. We also find that edits to $\text{W}_{a}^{(\ell)}$ have poor generalization (PS) in early layers, whereas high PS can be achieved at early layers by modifying either $\text{W}_{g}^{(\ell)}$ or $\text{W}_{o}^{(\ell)}$, consistent with their higher indirect effects as seen in Figure 5a.
@@
Where is the right place to apply ROME on Mamba? Figure 3 could suggest $\text{W}_{g}^{(\ell)}$, since the causal effect of $g_{i}$ states is mostly concentrated at the subject last token, similar to the behavior of MLPs in transformers (Meng et al., 2022a). Consistent with this is the architectural fact that, just as transformers’ $\text{W}_{down}^{(\ell)}$ connects to attention modules only through the residual stream, the output of $\text{W}_{g}^{(\ell)}$ does not flow through the Conv + SSM module—a module that other work has suggested might play role similar to that played by attention heads in transformers (Grazzi et al., 2024). And, indeed, we find that ROME can successfully insert facts by modifying $\text{W}_{g}^{(\ell)}$. On the other hand Figure 6a reveals sudden drops in efficacy and generalization at middle layer gates, suggesting that $\text{W}_{g}^{(\ell)}$ may be an unreliable mediator at some layers. Our experiments further show that the best performance for ROME is empirically achieved by modifying $\text{W}_{o}^{(\ell)}$. This is consistent with the fact that $o_{i}$ states show a stronger causal effect at the subject last token than $g_{i}$ states do (see Figure 3a). Additionally, ROME achieves better generalization (PS), competitive specificity (NS), and an overall better score (S) with $\text{W}_{o}^{(\ell)}$. We hypothesize that the strong performance of $\text{W}_{o}^{(\ell)}$ may be due to the the separation of roles between early-mid and later layers observed in Figures 2(b), 3a, and 5. Also see Appendix B where we isolate the contribution of $\text{W}_{o}^{(\ell)}$ by subtracting $\text{IE}_{s_{i}^{(\ell)}}+\text{IE}_{g_{i}^{(\ell)}}$ from $\text{IE}_{o_{i}^{(\ell)}}$, which reveal a critical role of $\text{W}_{o}^{(\ell)}$ in early-mid layers at subject last token position while mediating a fact.

### v2

@@ Abstract
We investigate the mechanisms of factual recall in the Mamba state space model. Our work is inspired by previous findings in autoregressive transformer language models suggesting that their knowledge recall is localized to particular modules at specific token locations; we therefore ask whether factual recall in Mamba can be similarly localized. To investigate this, we conduct four lines of experiments on Mamba. First, we apply causal tracing or interchange interventions to localize key components inside Mamba that are responsible for recalling facts, revealing that specific components within middle layers show strong causal effects at the last token of the subject, while the causal effect of intervening on later layers is most pronounced at the last token of the prompt, matching previous findings on autoregressive transformers. Second, we show that rank-one model editing methods can successfully insert facts at specific locations, again resembling findings on transformer LMs. Third, we examine the linearity of Mamba’s representations of factual relations. Finally we adapt attention-knockout techniques to Mamba in order to dissect information flow during factual recall. We compare Mamba directly to a similar-sized autoregressive transformer LM and conclude that despite significant differences in architectural approach, when it comes to factual recall, the two architectures share many similarities.
@@ 1 Introduction
Studies of autoregressive transformer language models’ (LMs) processing of factual statements such as The Eiffel Tower is located in Paris, have identified a localized pattern of internal computations when recalling facts (Meng et al., 2022a; b; Geva et al., 2023; Hernandez et al., 2023; Nanda et al., 2023), and have further found that those LMs can be edited by making single-layer rank-one changes in model parameters to alter a specific fact. Although these localized phenomena appear to generalize across autoregressive transformer LMs, the extent to which similar locality might appear in very different architectures—such as recurrent networks (RNNs)—has not yet been investigated.
@@
Our paper is a case study confronting a key methodological challenge that broadly faces interpretability researchers: as state-of-the-art neural network architectures evolve, we must ask, can the detailed analytical methods and tools developed for one neural architecture, such as transformer LMs, be generalized and applied to a different neural architecture, such as Mamba? In this paper we are able to answer the question with a qualified “yes”: we find that many of the methods used to analyze transformers can also provide insights on Mamba. We also discuss mismatches—that is, interpretation methods (such as path-dependent attention patching) that do not transfer to Mamba as easily due to architectural constraints.
@@
Finally, we apply methods for understanding the overall information flows in Mamba. Inspired by the findings of Hernandez et al. (2023), we measure the linearity of the relations between subject and object embeddings. And inspired by Geva et al. (2023), we examine information flow by adapting attention-blocking methods to the attention-free Mamba architecture.
@@ 2 Background on Mamba
Mamba, introduced in Gu & Dao (2023), is a recent family of language models based on state space models (SSMs). SSMs are designed to model the evolution of a hidden state across time with a first-order differential equation (Koopman et al., 1999; Durbin & Koopman, 2012), and when they are used as the recurrent state of an RNN, they can enable highly efficient parallelized training (Gu et al., 2021). To achieve good performance in language modeling, the Mamba SSM introduces input-dependent parameterization or selective-SSM instead of the traditional time-invariant SSMs. Mamba uses a special architecture called MambaBlock [^1] In their paper, Gu & Dao (2023) call this component Mamba—the same name as the LM family., which is stacked homogeneously, replacing both attention and MLP blocks used in transformer layers. Here, we focus on the different operations performed inside a MambaBlock.
@@ 3 Locating Key States for Factual Recall
Figure 2: (a) Activation patching. A state from the clean run $G$ is patched into its corresponding position in the corrupted run $G^{*}$. This has a downstream effect of potentially changing all the states that depend on the patched state in $G^{*}[\leftarrow h_{i}^{(\ell)}]$. (b) Average indirect effect of applying causal tracing on residual stream states ($h_{i}^{(\ell)}$ in Figure 1) across 400 different facts from the Relations dataset (see Section A.2).
@@
This subject-swapping approach follows the recommendation of Zhang & Nanda (2023) and has the advantage of using natural text perturbations to avoid introducing out-of-domain states to the model’s computation, as may happen when corrupting $s$ embeddings with Gaussian noise (the method used in Meng et al. (2022a)).
@@
Let $p(o)$, $p^{*}(o)$, and $p^{*}[{\leftarrow h_{i}^{(\ell)}}](o)$ denote the probability assigned to the correct answer $o$ in $G$, $G^{*}$, and $G^{*}[{\leftarrow h_{i}^{(\ell)}}]$ respectively. To measure the contribution of $h_{i}^{(\ell)}$ in recalling the fact $(s,r,o)$, we define its indirect effect (IE) as:
@@
In Figure 2(b) we plot the average indirect effect of restoring the residual states $h_{i}^{(\ell)}$ across different layer-token positions over 400 facts from the Relations dataset (Hernandez et al., 2023). The high IE observed at the late site (later layers at the last token) position is natural, as restoring a clean $h_{i}^{(\ell)}$ there will restore most of the model computation from $G$. However, Mamba also shows high causality at the early site (early-middle layers at the last subject token position). This is consistent with what Meng et al. (2022a) observed in the GPT family of language models.
@@
Figure 3: Average indirect effect of different states $o_{i}^{(\ell)}$, $g_{i}^{(\ell)}$, and $s_{i}^{(\ell)}$ over 400 facts from the Relations dataset (see Section A.2). For each layer $\ell$, states for a window of 10 layers around $\ell$ are restored from the clean run $G$.
@@
Figure 4: To probe for path-specific effects, (a) $h_{i}^{(\ell)}$ is restored from the clean run $G$ as in Figure 2(a). (b) Then, to reveal the role of the Conv + SSM contributions, $s_{i}$ states from the corrupted run $G^{*}$ are also patched to block the contributions from those paths.
@@
In Figure 3 we plot the average IE for $o_{i}^{(\ell)}$, $g_{i}^{(\ell)}$, and $s_{i}^{(\ell)}$. The plot for $o_{i}^{(\ell)}$ (Figure 3a) looks very similar to Figure 2(b), confirming that the output from MambaBlock has strong causal effects at both early and late sites. Interestingly, Figure 3c shows that the selective-SSM outputs $s_{i}^{(\ell)}$ have high IE only at the late site, resembling the behavior of attention modules in GPT models (Meng et al., 2022a). However, there is no state that appears to do the opposite; in other words, there is no state with strong effects at the early site and not at the late site (The gate output $g_{i}^{(\ell)}$ does have stronger IE at the early site, but these effects are very weak). To compare with autoregressive transformer LMs, activation patching results for Pythia-2.8b is shown in Figure 10 in Appendix D. This comparison reveals a key way how Mamba differs from transformers: while transformer MLP outputs have effects in the early site and not the late site, in Mamba there is no similar state that specializes only at the early site, at which factual recall would be expected to occur. This presents the question: which parameters in Mamba mediate factual recall?
@@
To investigate this question, we replicate an experiment from Meng et al. (2022a) to probe path-specific effects (Pearl, 2022) by severing a path from the causal graph and monitoring its effect. Here, we are interested in understanding the effect of the contributions from $g_{i}$, $s_{i}$, and $o_{i}$ (i.e. states that are processed by $\text{W}_{g}$, Conv + SSM, and $\text{W}_{o}$ respectively) while recalling a fact. First, in the corrupted run $G^{*}$, at token position $i$, we cache all the contributions from the $s_{i}$ paths as $s_{i}^{*}=\{s_{i}^{*(\ell)}|\;\ell\in[1,L]\}$. Then in the patched run $G^{*}[{\leftarrow h_{i}^{(\ell)}}]$, we restore $h_{i}^{(\ell)}$ that was cached from the clean run $G$ into its corresponding state (as in Figure 2(a)), but with an additional modification: to understand the contribution from the $s_{i}$ paths, we sever those paths by also patching $s_{i}^{*}$ (cached from the corrupted run $G^{*}$) to their corresponding locations (see Figure 4). The same experiment is replicated to understand the contributions of $g_{i}$ and $o_{i}$ states. We note that severing the $o_{i}^{(\ell)}$ will sever $s_{i}^{(\ell)}$ and $g_{i}^{(\ell)}$ as well (see Figure 1).
@@
In Figure 5 we plot the average results of this experiment for token positions (a) $i=\textit{subject last}$ and (b) $i=\textit{prompt last}$ over 400 examples randomly sampled from the Relations dataset. The key findings can be understood by examining the gap between the purple bars and the green, red, and blue bars: a large gap indicates a strong mediating role for Conv + SSM, $\text{W}_{g}$, or $\text{W}_{o}$ parameters, respectively. At the early site at the subject last token, both the Conv + SSM and $\text{W}_{g}$ have a strong role, but $\text{W}_{o}$ plays an even larger role than either. Yet the strongest mediator at the late site is also $\text{W}_{o}$. This experiment highlights the importance of $\text{W}_{o}$ in both stages of predicting a fact. But it also suggests that Mamba does not separate early-site factual recall between these groups of parameters as cleanly as transformers. However, Figure 5 reveals a clean separation of roles between early to mid and later layers, analogous to the findings of Hernandez et al. (2023) in transformer LMs. We also note that this division of responsibilities between layers can be more sharply noticed in Mamba when compared to transformers LMs (compare Figure 5 with Figure 11).
@@ 4 Editing Facts With ROME
Having begun to characterize the locations of important states for factual recall, we now investigate whether factual recall behavior can be edited. In particular, we apply the ROME (Rank One Model Editing, Meng et al., 2022a) technique to Mamba. ROME begins with the observation that any linear transformation can be considered as an associative memory (Anderson, 1972; Kohonen, 1972), mapping a set of keys $K=[k_{1}|k_{2}|\dots]$ to their corresponding values $V=[v_{1}|v_{2}|\dots]$, and uses this to edit factual associations in transformer LMs. Here, we apply the technique to a particular set of linear transformations within Mamba, and report our editing success on each. [^2] Further motivating these experiments, previous work has shown that the locations identified by activation patching techniques are not necessarily those which have the strongest edit performance  (Hase et al., 2024).
[…]
@@ 4.1 Applying ROME in Mamba
We apply ROME on the three different projection matrices of Mamba: $\text{W}_{a}^{(\ell)}$ which affects only the Conv + SSM path, $\text{W}_{g}^{(\ell)}$ which affects only the gating path, and $\text{W}_{o}^{(\ell)}$, the final output of the MambaBlock, which is added to the residual state. We plot ROME performance on different projection matrices ($\text{W}_{a}^{(\ell)}$, $\text{W}_{g}^{(\ell)}$, and $\text{W}_{o}^{(\ell)}$) across all the layers in Figure 6a.
@@
To evaluate editing performance, we use the CounterFact dataset from Meng et al. (2022a). CounterFact contains 20K counterfactual examples in the form $(s,r,o\rightarrow o^{*})$, where $o$ is the correct answer to the prompt $x=(s,r)$, and $o^{*}$ is the object which is to be inserted as the new answer to the prompt (See Section A.1 for details). We select the first 2000 examples from this dataset for our module-layer sweep. We use the original evaluation matrices in Meng et al. (2022a) to measure ROME edit performance.
@@
    Generalization (PS): A successful edit should be persistent across different paraphrases of $(s,r)$. For each of the request instances $(s,r,o\rightarrow o^{*})$, $p(o^{*})>p(o)$ is checked post-edit with a set of different rephrasings $x_{p}\sim\mathcal{P}_{r}(s)$ of the prompt $x=(s,r)$, where $\mathcal{P}_{r}$ denotes a set of paraphrased templates for the relation $r$.
@@
    Specificity (NS): Finally, the edit should be specific to $\mathcal{P}_{r}(s)$ and should not additionally change the mapping of some nearby subject $s_{n}$ to $o^{*}$. To evaluate the specificity of an edit we measure $p(o_{n})>p(o^{*})$ with $\mathcal{P}_{r}(s_{n})$ for a set of nearby factual associations $\{(s_{n},r,o_{n})\,|\,o_{n}\neq o^{*}\}$.
@@
Figure 6a shows that ROME can achieve high scores (S) for a range of early to middle layers by modifying any one of the projection matrices $\text{W}_{a}^{(\ell)}$, $\text{W}_{g}^{(\ell)}$, or $\text{W}_{o}^{(\ell)}$, matching observations made by Hase et al. (2024) regarding transformer LMs. However, we found that performance does depend on the location of the edit. For example, in the case of $\text{W}_{g}^{(\ell)}$ and $\text{W}_{o}^{(\ell)}$, the score (S) and generalization (PS) drops after around layer 43. This is consistent with our findings from the path-blocking experiment in Figure 5a. We also find that edits to $\text{W}_{a}^{(\ell)}$ have poor generalization (PS) in early layers, whereas high PS can be achieved at early layers by modifying either $\text{W}_{g}^{(\ell)}$ or $\text{W}_{o}^{(\ell)}$, consistent with their higher indirect effects as seen in Figure 5a.
@@
Where is the right place to apply ROME on Mamba? Figure 3 could suggest $\text{W}_{g}^{(\ell)}$, since the causal effect of $g_{i}$ states is mostly concentrated at the subject last token, similar to the behavior of MLPs in transformers (Meng et al., 2022a). Consistent with this is the architectural fact that, just as transformers’ $\text{W}_{down}^{(\ell)}$ connects to attention modules only through the residual stream, the output of $\text{W}_{g}^{(\ell)}$ does not flow through the Conv + SSM module—a module that other work has suggested might play a role similar to that played by attention heads in transformers (Grazzi et al., 2024). And, indeed, we find that ROME can successfully insert facts by modifying $\text{W}_{g}^{(\ell)}$. On the other hand Figure 6a reveals sudden drops in efficacy and generalization at middle layer gates, suggesting that $\text{W}_{g}^{(\ell)}$ may be an unreliable mediator at some layers. Our experiments further show that the best performance for ROME is empirically achieved by modifying $\text{W}_{o}^{(\ell)}$. This is consistent with the fact that $o_{i}$ states show a stronger causal effect at the subject last token than $g_{i}$ states do (see Figures 3a and 5a). Additionally, ROME achieves better generalization (PS), competitive specificity (NS), and an overall better score (S) with $\text{W}_{o}^{(\ell)}$. We hypothesize that the strong performance of $\text{W}_{o}^{(\ell)}$ may be due to the the separation of roles between early-mid and later layers observed in Figures 2(b), 3a, and 5. Also see Appendix C where we isolate the contribution of $\text{W}_{o}^{(\ell)}$ by subtracting $\text{IE}_{s_{i}^{(\ell)}}+\text{IE}_{g_{i}^{(\ell)}}$ from $\text{IE}_{o_{i}^{(\ell)}}$, which reveal a critical role of $\text{W}_{o}^{(\ell)}$ in early-mid layers at subject last token position while mediating a fact.